OpenStudy (anonymous):

How much water must be added to 12 L of a 40% solution of alcohol to obtain a 30% solution? A. .5 L B. 5 L C. .4 L D. 4 L

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you have 12 L of a 40% alcohol solution how much pure alcohol do you have?

OpenStudy (anonymous):

C?

jimthompson5910 (jim_thompson5910):

40% of 12 = ??

OpenStudy (anonymous):

wait sorry i waking thinking about making the % in to a decimal

jimthompson5910 (jim_thompson5910):

40% of 12 = 0.4*12 = ??

OpenStudy (anonymous):

its 4.8

jimthompson5910 (jim_thompson5910):

ok so you have 4.8 L of pure alcohol

jimthompson5910 (jim_thompson5910):

let x = amount of water you want to add to the solution

jimthompson5910 (jim_thompson5910):

your initial solution is 12 L when you add x liters of water, you jump to 12+x liters total (of mixed solution)

jimthompson5910 (jim_thompson5910):

the pure alcohol amount stays the same...so you still have 4.8 L of pure alcohol because you aren't adding/subtracting any alcohol

OpenStudy (anonymous):

so were trying to take away pure alcholo right ? to be 30 %

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

you don't add or remove any amount of alcohol

jimthompson5910 (jim_thompson5910):

you just add more water to dilute the solution to make the alcohol percentage lower

OpenStudy (anonymous):

and how do we do that ?

jimthompson5910 (jim_thompson5910):

what this means is that [ (amount of pure alcohol)/(total amount) ]*100 = percentage of alcohol in solution [ (4.8)/(12+x) ] * 100 = 30 (4.8)/(12+x) = 30/100 (4.8)/(12+x) = 0.3 4.8 = 0.3(12+x) 4.8 = 3.6+0.3x 4.8-3.6 = 0.3x 1.2 = 0.3x 0.3x = 1.2 x = 1.2/0.3 x = 4 So you need to add 4 liters of water to turn 12 L of a 40% solution into a 30% solution